# ThinkQE: Query Expansion via an Evolving Thinking Process

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have led to strong performance in query expansion Gao et al. (2022); Wang et al. (2023); Jagerman et al. (2023); Mackie et al. (2023); Shen et al. (2024); Lei et al. (2024), particularly due to their ability to rapidly adapt to new domains without requiring additional training. However, existing LLM-based methods often pay limited attention to exploration and result diversity. As illustrated in Table 1, we observe that current approaches – such as HyDE – tend to generate overly confident expansions that focus narrowly on a single interpretation of the input query. This behavior can be attributed to the model’s reliance on its internal knowledge and high-probability completions Yona et al. (2024); Ohi et al. (2024); Sun et al. (2025), which may suppress alternative formulations or less common aspects of the query. This lack of breadth limits the retrieval of documents reflecting alternative scenarios or requiring more nuanced reasoning.
@@ 2 Method
We introduce ThinkQE, a query expansion framework that tightly integrates LLM-based thinking process with evolving corpus interaction. The method is designed to enable exploration of the query space through thinking processes and evolving refinement based on retrieval feedback from the corpus. The overall process proceeds in multiple rounds. At each round, an LLM performs thinking-augmented expansion based on the original query and newly retrieved documents from the corpus, which in turn informs subsequent retrieval and expansion steps. This section describes each component of the method in detail.
@@ 2.2 Expansion via Thinking Process
Given a question "{$q$}" and its possible answering passages (most of these passages are wrong) enumerated as:
@@
[…]
@@ 2.3 Evolution via Corpus Feedback
2. Redundancy Filtering: To promote diversity and avoid repetition, we exclude documents that (a) appear in the blacklist $\mathcal{B}_{t}$, or (b) were among the top-$K$ results in the previous round $\mathcal{D}_{t-1}$. We then select the top-$K$ documents from the remaining candidates: $\mathcal{D}_{t}^{\text{new}}=\text{TopK}(\mathcal{R}_{t}\setminus(\mathcal{B}_{t}\cup\mathcal{D}_{t-1})).$ The blacklist is updated to include all documents that were filtered out in this round.
@@
This loop can be repeated for any number of rounds $T$, depending on resource constraints or desired depth.
[…]
@@ 3.1 Setup
We use the QWEN-R1-Distill-14B model (DeepSeek-AI, 2025) to generate thinking-based query expansions, sampling outputs with a temperature of 0.7. The BM25 retrieval is performed using Pyserini Lin et al. (2021) with default hyperparameters. At each round, ThinkQE uses the top-5 retrieved documents (truncated to 128 tokens for DL benchmarks and 512 tokens for BRIGHT) to prompt the LLM, and samples 2 candidate expansions to enhance diversity. We set the total number of interaction rounds to 3, for a balance between efficiency and effectiveness.
@@
ThinkQE-14B 45.9 68.8 89.3 43.9 64.7 87.8
@@
HyDE-14B Zero-shot 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
LameR-14B Zero-shot 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
ThinkQE-14B (Ours) Zero-shot 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
Table 4: Results on the StackExchange domain of the BRIGHT benchmark in terms of nDCG@10. The best and the second best results across all models are in bold and underlined font, respectively. BM25+GPT-4o-CoT refers to applying BM25 to queries rewritten by GPT-4o with CoT reasoning traces included. All models except BM25+GPT-4o-CoT are performed on the original query. *Rank-K-32B performs computationally expensive listwise reranking over the top-20 documents.
@@
On the BRIGHT benchmark, we consider three categories of baselines: (1) LLM-based embedding models such as GritLM-7B (Muennighoff et al., 2025), GTE-Qwen-7B (Li et al., 2023), and ReasonIR-8B Shao et al. (2025), all trained on massive amounts of retrieval data; (2) LLM-based rerankers, including RankGPT4 (zero-shot) (Sun et al., 2023), RankZephyr-7B (distilled from GPT-4) (Pradeep et al., 2023), Rank1-14B (distilled from DeepSeek-R1-685B) (Weller et al., 2025), Rank-R1-14B (trained via reinforcement learning) (Zhuang et al., 2025), and Rank-K-32B (distilled from DeepSeek-R1-685B) Yang et al. (2025). Rank1-14B, Rank-R1-14B, and Rank-K-32B explicitly incorporate a thinking process during reranking; and (3) Query expansion methods such as HyDE and LameR, use the same underlying model as ThinkQE but do not incorporate any explicit thinking process. [^3] We provide a detailed analysis of the no-thinking setting for fair comparison with ThinkQE in Section 4.1. Our method ThinkQE is evaluated in a zero-shot configuration across all datasets.
@@ 3.2 Main Results
On the BRIGHT benchmark, ThinkQE remains the strongest among zero-shot query expansion methods, achieving an average nDCG@10 of 34.9. While Rank-K-32B achieves the highest overall score (37.9), it relies on R1 distillation and listwise reranking over GPT-4o-augmented results, making it significantly more resource-intensive. In contrast, ThinkQE operates in a fully training-free setting and still outperforms several more expensive rerankers, including RankGPT4 (24.7) and Rank1-14B (31.7). Beyond its efficiency, ThinkQE delivers consistently strong performance across domains, ranking first or second in four sub-domains.
@@ A.1 Dataset Statistics
Table 7: Dataset Statistics
A.2 Detailed Results on the Impact of the Thinking Process
The detailed results across all domains on the impact of the thinking process are provided in Table 8.
@@ A.2 Detailed Results on the Impact of the Thinking Process
A.3 Detailed Results on the Impact of Evolving Corpus Interaction
The detailed results across all domains on the impact of the evolving corpus interaction are provided in Table 9.
@@ A.3 Detailed Results on the Impact of Evolving Corpus Interaction
A.4 Detailed Results on the Core Components of the Evolving Interaction Process
The detailed results across all domains on the impact of the expansion accumulation and redundancy filter mechanisms are provided in Table 10.
@@ A.4 Detailed Results on the Core Components of the Evolving Interaction Process
Table 10: Detailed results on the impact of the expansion accumulation and redundancy filter mechanisms.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have led to strong performance in query expansion Gao et al. (2022); Wang et al. (2023); Jagerman et al. (2023); Mackie et al. (2023); Shen et al. (2024); Lei et al. (2024), particularly due to their ability to rapidly adapt to new domains without requiring additional training. However, existing LLM-based methods often pay limited attention to exploration and result diversity. As illustrated in Table 1, we observe that current approaches, such as HyDE, tend to generate overly confident expansions that focus narrowly on a single interpretation of the input query. This behavior can be attributed to the model’s reliance on its internal knowledge and high-probability completions Yona et al. (2024); Ohi et al. (2024); Sun et al. (2025), which may suppress alternative formulations or less common aspects of the query. This lack of breadth limits the retrieval of documents reflecting alternative scenarios or requiring more nuanced reasoning.
@@ 2 Method
We introduce ThinkQE, a query expansion framework that tightly integrates LLM-based thinking process with evolving corpus interaction. ThinkQE follows prior work in generating query expansions using retrieved documents but distinguishes itself through its design of thinking augmentation and iterative corpus interactions. The method is designed to enable exploration of the query space through thinking processes and evolving refinement based on retrieval feedback from the corpus. The overall process proceeds in multiple rounds. At each round, an LLM performs thinking-augmented expansion based on the original query and newly retrieved documents from the corpus, which in turn informs subsequent retrieval and expansion steps. The following subsections describe each component of the method in detail.
@@ 2.2 Expansion via Thinking Process
Given a question "{qq}" and its possible answering passages (most of these passages are wrong) enumerated as:
@@
Table 2: Prompt used in ThinkQE for the thinking-based expansion process. $\{\cdot\}$ denotes the placeholder for the corresponding query and top-K documents.
@@ 2.3 Evolution via Corpus Feedback
2. Redundancy Filtering: To promote diversity and avoid repetition, we exclude documents that (a) appear in the blocklist $\mathcal{B}_{t}$, or (b) were among the top-$K$ results in the previous round $\mathcal{D}_{t-1}$. We then select the top-$K$ documents from the remaining candidates: $\mathcal{D}_{t}^{\text{new}}=\text{TopK}(\mathcal{R}_{t}\setminus(\mathcal{B}_{t}\cup\mathcal{D}_{t-1})).$ The blocklist is updated to include all documents that were filtered out in this round.
@@
This loop can be repeated for any number of rounds $T$, depending on resource constraints or desired depth. Notably, as the query grows longer, successive expansions may dilute or replace the original intent. To mitigate this, we follow Zhang et al. (2024) and repeat the original query $n$ times in the final reformulation, with $n=\frac{\text{len}(\text{expansions})}{\text{len}(q_{0})\times\lambda},\quad\lambda=3.$ Here, $\text{len}(\text{expansions})$ refers to the total word count of all expansion segments, and $\text{len}(q_{0})$ is the word count of the original query. This repetition reinforces the core semantics of the original query during iterative refinement.
@@ 3.1 Setup
We use the QWEN-R1-Distill-14B model (DeepSeek-AI, 2025) to generate thinking-based query expansions, sampling outputs with a temperature of 0.7. The BM25 retrieval is performed using Pyserini Lin et al. (2021) with default hyperparameters. At each round, ThinkQE uses the top-5 retrieved documents (truncated to 128 tokens for DL benchmarks and 512 tokens for BRIGHT) to prompt the LLM, and samples 2 candidate expansions to enhance diversity. We set the total number of interaction rounds to 3, for a balance between efficiency and effectiveness. Besides the QWEN-R1-Distill-14B model, we also evalute ThinkQE on a variety of reasoning models, including QWEN3-8B, QWEN3-14B Yang et al. (2025a), OpenThinker2-7B Meta (2023), and Phi-4-Reasoning-14B Abdin et al. (2025).
@@
CSQE 43.6 63.4 87.6 - - -
ThinkQE (ours)
w. R1-14B 45.9 68.8 89.3 43.9 64.7 87.8
w. QWEN3-8B 44.5 65.0 87.9 41.9 62.8 88.0
w. QWEN3-14B 45.2 64.9 88.4 42.4 63.5 88.4
w. OpenThinker2-7B 44.8 65.3 87.3 43.2 63.5 87.9
w. Phi4-Reasoning-14B 44.0 65.0 87.1 43.0 63.9 87.2
@@
HyDE-R1-14B Zero-shot 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
LameR-R1-14B Zero-shot 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
ThinkQE (ours)
   R1-14B Zero-shot 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
   QWEN3-8B Zero-shot 49.8 55.3 27.6 36.7 19.9 29.0 28.3 35.2
   QWEN3-14B Zero-shot 51.5 53.2 27.8 37.2 22.0 16.1 27.5 33.6
   OpenThinker2-7B Zero-shot 50.5 54.1 25.8 36.7 18.1 28.2 28.9 34.6
   Phi-4-Reasoning-14B Zero-shot 51.8 53.5 29.7 38.5 21.8 29.3 27.7 36.0
Table 4: Results on the StackExchange domain of the BRIGHT benchmark in terms of nDCG@10. The best and the second best results across all models are in bold and underlined font, respectively. All models are performed on the original query. BM25+GPT-4o-CoT refers to using BM25 retrieval results on queries rewritten by GPT-4o with chain-of-thought reasoning traces for reranking. *Rank-K-32B performs computationally expensive listwise reranking over the top-20 documents.
@@
On the BRIGHT benchmark, we consider three categories of baselines: (1) LLM-based embedding models such as GritLM-7B (Muennighoff et al., 2025), GTE-Qwen-7B (Li et al., 2023), and ReasonIR-8B Shao et al. (2025), all trained on massive amounts of retrieval data; (2) LLM-based rerankers, including RankGPT4 (zero-shot) (Sun et al., 2023), RankZephyr-7B (distilled from GPT-4) (Pradeep et al., 2023), Rank1-14B (distilled from DeepSeek-R1-685B) (Weller et al., 2025), Rank-R1-14B (trained via reinforcement learning) (Zhuang et al., 2025), and Rank-K-32B (distilled from DeepSeek-R1-685B) Yang et al. (2025b). Rank1-14B, Rank-R1-14B, and Rank-K-32B explicitly incorporate a thinking process during reranking; and (3) Query expansion methods such as HyDE and LameR, which use the same underlying model as ThinkQE but do not incorporate any explicit thinking process. [^3] We provide a detailed analysis of the no-thinking setting for fair comparison with ThinkQE in Section 4.1. Our method ThinkQE is evaluated in a zero-shot configuration across all datasets.
@@ 3.2 Main Results
On the BRIGHT benchmark, ThinkQE remains the strongest among zero-shot query expansion methods, achieving an average nDCG@10 of 36.0 using the Phi-4-Reasoning-14B model. While Rank-K-32B achieves the highest overall score (37.9), it relies on R1 distillation and listwise reranking over GPT-4o-augmented retrieval results, making it significantly more resource-intensive. In contrast, ThinkQE operates in a fully training-free setting and still outperforms several more expensive rerankers, including RankGPT4 (24.7) and Rank1-14B (31.7). Beyond its efficiency, ThinkQE delivers consistently strong performance across domains, ranking first or second in four sub-domains.
@@ A.1 Dataset Statistics
Table 7: Dataset Statistics.
A.2 Impact of the Thinking Process
Results across all domains on the impact of the thinking process are provided in Table 10.
A.3 Impact of Evolving Corpus Interaction
Results across all domains on the impact of the evolving corpus interaction are provided in Table 11.
A.4 Core Components of the Evolving Interaction Process
Results across all domains on the impact of the expansion accumulation and redundancy filter mechanisms are provided in Table 12.
A.5 Impact of $\lambda$
Table 13 presents detailed results analyzing the impact of $\lambda$, which influences the repetition frequency of the original query during expansion. The results demonstrate that performance differences are small when varying $\lambda$ from 3 to 6. However, lower $\lambda$ values tend to cause excessive repetition of the original query, which generally hurts performance.
A.6 Effectiveness-Efficiency Analysis on ThinkQE
We provide latency and performance trade-off results comparing model size scaling, the thinking process, and multiple rounds in Table 14, using the DeepSeek-R1-Distill-Qwen-14B and DeepSeek-R1-Distill-Qwen-32B models evaluated on a single H100 GPU. Our results show that involving the thinking process and multiple rounds increases latency. However, scaling model size alone yields limited improvements relative to latency increase: moving from R1-14B without thinking (3.71 s/query, 29.8) to R1-32B without thinking (7.88 s/query) improves Bright Avg. by just +0.6 points (30.4) while more than doubling latency. Adding the thinking process, although it increases latency more, is substantially more effective: applying 1 round of thinking to R1-14B boosts performance by +2.7 points (32.5) compared to R1-14B without thinking, with latency rising to 15.40 s/query. Even with thinking enabled, scaling from R1-14B to R1-32B brings only a small additional gain (+0.4 points). Meanwhile, multi-round corpus interaction offers a more efficient path to higher effectiveness, with R1-14B 3-round reaching 34.9 (+2.4 over its 1-round version) and outperforming R1-32B 1-round.
A.7 ThinkQE on Non-Web Search Datasets
We evaluate ThinkQE on two additional non-web search datasets, TREC-Covid and Scifact. The results in Table 8 show that ThinkQE remains highly competitive and often outperforms the baselines. Note that both MILL and LameR are based on the powerful, closed-source GPT-3.5-Turbo model.
Method TREC-Covid Scifact
BM25 59.5 67.9
Contriever (FT) 59.6 67.7
HyDE 59.3 69.1
Query2Doc 72.2 68.6
MILL 75.3 71.4
LameR 75.8 73.5
CSQE 74.2 69.6
ThinkQE 76.1 73.3
Table 8: Results (NDCG@10) on non-web search datasets.
A.8 Significance Test Results of ThinkQE
We conduct significance testing by comparing ThinkQE with the two most relevant baselines we reimplemented on BRIGHT: HyDE and LameR using the same QWEN-R1-Distill-14B model. The results are presented in Table 15. Significance tests were performed using a t-test with a p-value threshold of 0.05. The results show that ThinkQE is significantly better than HyDE and LameR on 6 out of 7 and 5 out of 7 domains, respectively, demonstrating its effectiveness.
A.9 Impact of Number of Rounds on Domain-Specific Dataset
To further analyze the potential topic drift issue during the iterative process, we examine performance changes across one to three rounds using the domain-specific TREC-Covid dataset. The results, presented in Table 9, demonstrate that increasing from one to two rounds improves the performance, while extending to three rounds largely maintains performance.
Round NDCG@10
1 75.2
2 76.2
3 76.1
Table 9: Performance across iterative rounds on the domain-specific TREC-COVID dataset.
@@ A.9 Impact of Number of Rounds on Domain-Specific Dataset
Table 12: Detailed results on the impact of the expansion accumulation and redundancy filter mechanism.
$\mathbf{\lambda}$ Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
1 38.2 47.5 25.0 32.1 18.4 25.5 24.0 30.1
2 44.9 50.1 28.2 37.0 19.3 27.0 26.8 33.3
3 46.7 51.2 29.4 39.3 20.8 28.3 28.4 34.9
4 48.9 52.3 28.2 38.4 20.2 28.1 29.4 35.1
5 49.7 52.5 29.5 37.7 19.6 28.9 27.6 35.1
6 49.1 51.8 29.5 39.1 19.4 29.0 28.8 35.2
Table 13: Detailed results on the impact of $\lambda$.
Model Thinking Round Latency (second/query) Bright Avg.
R1-14B No 1 3.71 29.8
R1-32B No 1 7.88 30.4
R1-14B Yes 1 15.40 32.5
R1-32B Yes 1 30.53 32.9
R1-14B Yes 3 45.44 34.9
Table 14: Effectiveness-Efficiency Analysis on ThinkQE.
Method Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
HyDE 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
LameR 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
ThinkQE $47.3^{\dagger\ddagger}$ $52.5^{\dagger\ddagger}$ $29.2^{\dagger\ddagger}$ $40.0^{\dagger\ddagger}$ 19.3 $28.0^{\dagger}$ $27.9^{\dagger\ddagger}$ 34.9
Table 15: Significance testing results. $\mathbf{\dagger}$ and $\mathbf{\ddagger}$ mean ThinkQE performs significantly better than HyDE and LameR, respectively, as determined by a t-test with p-value 0.05 as threshold.